Activity
From 03/14/2025 to 04/12/2025
03/14/2025
- 11:45 AM Task #1086 (New): Reboot for RPi
- 11:44 AM SBS DAQ GEp GEMS Task #1053 (Closed): GEM LV current readback
- 11:43 AM SBS DAQ GEp CDet Task #1059 (In Progress): VETROC CDet software
- 11:42 AM SBS DAQ GEp CDet Task #1060 (In Progress): Calibration scripts for CDet
- 11:40 AM SBS DAQ GEp GEMS Bug #1068 (Resolved): swap optical receivers
- 11:40 AM SBS DAQ GEp ECal Task #1071 (In Progress): Software for calibration and set threshold for ECAL
- 11:39 AM SBS DAQ GEp GEMS Task #1081 (In Progress): Modify readout list to reload firmware at prestart
Also available in: Atom